# How to Understand Whole Software Repository?

### v1

@@ 4.4.2. Hyper Parameter Analysis
We further analyzed the impact of the iterations number in MCTS. We set the maximum number of iterations to 50, 200, and 600, and limited the maximum iteration time to 300 seconds. The results are shown in Table 5. We found that: (1) As the number of iterations increases, RepoUnderstander solves more actual issues. This shows that as the number of iterations rounds increases, agents will collect more repository information, i.e., they will have more experience with the repository, resulting in a higher problem solving rate; (2) As the number of iterations increases, we found that the relative improvement in problem solving gradually decreases. Specifically, the improvement of 50 iterations is significant compared to no iterations, but the relative improvement of the subsequent 200 and 600 iterations decreases. This may be because in the early stage, agents can quickly search and summarize relevant experience, but as the number of iterations increases, the convergence speed of the model gradually slows down, and the contribution of new information to performance improvement becomes smaller; (3) We observed that as the number of iterations increases from 200 to 600, the apply rate decreases. This phenomenon indicates that as the number of iterations increases, the model may be affected by some interference information when generating results, resulting in a decrease in the quality of the generated results. Therefore, when selecting the number of iterations, it is necessary to consider avoiding the influence of excessive interference information.
Iters Resolved Apply
@@
Table 5. Hyperparameter results of RepoUnderstander.
Figure 5. Distribution of results in SWE-bench Lite.
4.5. Case Study
4.5.1. Wrong Reason
Although RepoUnderstander achieved better results than other methods in the ASE task, the task is still challenging (even when RepoUnderstander and SWE-agent worked together, only 80/300 instances were fully automatically solved). To guide the optimization of subsequent work, we analyzed the specific reasons for the unsolved issues and divided the failure types into three categories: wrong location, apply patch failed, and wrong patch. Wrong location means that the agent did not locate the root cause of the problem and mistakenly modified the code in other correct locations. Apply patch failed means that the agent did not generate a syntactically correct patch and could not be directly applied to the existing version of the repository. Wrong patch means that when the bug location and patch syntax are correct, the repaired code cannot completely solve the problem, i.e., the test case does not pass completely.
We compared the results of RepoUnderstander and the current SOTA agent SWE-agent on SWE-bench Lite, as shown in the figure. The results show that: (1) In terms of bug location, our method generated patches in the wrong location in 45.0% of the tasks, while SWE-agent generated patches in 62.0%. This shows that our method performs better in correctly locating bugs, and the global experience at the repository level plays a key role in fault location. (2) In terms of patch applicability, our method did not generate patches in 13.7% of the tasks, while SWE-agent generated 5.0%. (3) In terms of patch correctness, our method generated incorrect patches in 20.0% of the tasks, while SWE-agent generated 15.0%. These indicators show that our method is more accurate in locating bug locations, but has certain weaknesses in generating correct patches and ensuring patch applicability. This may be because the execution feedback module of SWE-agent allows the agent to iteratively generate and verify the correctness of patches. This also shows the importance of execution feedback for solving practical problems. Therefore, by combining the global experience at the repository level and an effective execution feedback mechanism, we can further optimize the effect of automated software repair.
4.5.2. Resolved Tasks Study
Figure 6. Case of django-15498. (Partial wrokflow of RepoUnderstander and SWE-agent)
Figure 7. Case of matplotlib-26011. (Partial wrokflow of RepoUnderstander and SWE-agent)
To further analyze the performance of RepoUnderstander in real-world problems, we selected two examples from SWE-bench (as shown in Figures 6 and 7) and compared them with the SOTA solution, SWE-agent. By analyzing the oracle patch, we determined the actual fault location that needs to be modified. Among them, the green highlight indicates that the agents correctly tracked the target that needs to be modified, and the yellow highlight indicates that the agents incorrectly identified the target that is not to be modified.
As shown in Figure 6, RepoUnderstander correctly explored the issue-related function to be modified was_modified_since in the MCTS-based Repository Exploration stage; in the subsequent summary stage, it clearly pointed out that the function needed to be updated; finally, under the guidance of global experience, the agents accurately searched and located the was_modified_since function in the dynamic information acquisition stage and the patch generation stage, and successfully implemented the correct modification. In contrast, in SWE-agent, agents can search for files and view file contents in a fixed window size through computer interface operations. However, due to the lack of guidance from repositroy experience knowledge, SWE-agent mistakenly located the function get_conditional_response in the local search. Although this function has a certain relevance to solving the issue, it is not directly related. Therefore, modifying the wrong place resulted in the failure to solve the task.
In Figure 7, we analyzed and found that solving the matplotlib-26011 task can be achieved by modifying the _set_lim or set_xlim function. Both RepoUnderstander and SWE-agent correctly located the position to be modified. However, due to the running feedback and iteration capabilities of SWE-agent, the agents finally generated the correct patch through feedback information. This shows the importance of execution feedback for solving practical problems. Since RepoUnderstander and SWE-agent have certain complementarity, in the future we will combine global experience and effective execution feedback mechanism to further optimize the effect of automated software repair.
@@ 5. Limitation
5.1. Resource Overhead.
Although RepoUnderstander aims to guide large language models to fully understand the whole software repository to effectively solve the challenges in automatic software engineering (ASE), the Monte Carlo Tree Search (MCTS) process does require a certain amount of resource consumption. Specifically, we set the maximum number of iterations to 600 and the maximum search time to 300 seconds to ensure that the model can fully explore the search space and accurately evaluate the rewards of different paths. However, such settings are controllable and adjustable to adapt to different application scenarios and resource constraints. Through reasonable parameter adjustment, the best balance between resource consumption and result accuracy can be found. In addition, as shown in Table 5, only 50 iterations can also achieve results that are superior to other agents. Further research may discover more efficient strategies to reduce resource requirements while maintaining or improving agents performance.
5.2. Runtime feedback.
RepoUnderstander aims to study how to understand the entire software repository, and has emerged effective ASE capabilities through modules such as software knowledge graph construction and MCTS-enhanced repository understanding, especially in tasks such as fault location that rely on effective understanding of the entire repository knowledge. SWE-agent (Yang et al., [n. d.]) uses information such as execution feedback to assist in verifying the correctness of generated patches, and iteratively repairs patches to pass tests, which has a certain effect on improving the correctness of patch generation. In Table 2, we analyze the distribution of problem solving by RepoUnderstander and SWE-agent, and find that the two methods have a certain complementarity, jointly solving 26.67% of the problems on SWE-bench Lite. Inspired by this, in the future, further combining RepoUnderstander with runtime feedback is an effective way to enhance ASE capabilities.
@@ 6. Conclusion
This paper highlights the significance of understanding the whole software repository as a critical path to achieving automatic software engineering (ASE). To this end, we present a novel LLM-based agent method named RepoUnderstander, which guides agents to comprehensively understand entire repositories. Concretely, RepoUnderstander first constructs a repository knowledge graph to condense extensive and complex repository-level information into a hierarchical structure. Subsequently, we enhance the agents’ repository understanding through a Monte Carlo Tree Search (MCTS) enhanced repository exploration strategy. Finally, we guide the agents to summarize, analyze, and plan according to global experiences. Then, they can manipulate the tools to dynamically acquire information and generate patches to solve real-world GitHub issues.
We demonstrate the superior performance of RepoUnderstander through extensive experiments and comprehensive analyses. Our method achieves state-of-the-art performance on the SWE-bench Lite benchmark, outperforming existing RAG-based and agent-based systems. Furthermore, we address the issue of information asymmetry in the SWE-bench Lite dataset by proposing a FIX version that integrates new feature interface specifications into the problem description. Additionally, we conduct ablation studies to analyze the effectiveness of various components of RepoUnderstander and identify areas for further improvement. Our findings emphasize the importance of global repository experiences and the potential benefits of integrating runtime feedback mechanisms. Future work will focus on combining global experiences with runtime feedback mechanisms to further enhance the capabilities of LLM-based agents in solving complex software engineering tasks.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v2

@@ 4.4.3. Hyper Parameter Analysis
MCTS_Iters Resolved Apply
@@
Table 6. Hyperparameter results.
We further analyzed the impact of the iterations number in MCTS. We set the maximum number of iterations to 50, 200, and 600, and limited the maximum iteration time to 300 seconds. The results are shown in Table 6. We found that: (1) As the number of iterations increases, LingmaAgent solves more actual issues. This shows that as the number of iterations rounds increases, agents will collect more repository information, i.e., they will have more experience with the repository, resulting in a higher problem solving rate; (2) As the number of iterations increases, we found that the relative improvement in problem solving gradually decreases. Specifically, the improvement of 50 iterations is significant compared to no iterations, but the relative improvement of the subsequent 200 and 600 iterations decreases. This may be because in the early stage, agents can quickly search and summarize relevant experience, but as the number of iterations increases, the convergence speed of the model gradually slows down, and the contribution of new information to performance improvement becomes smaller; (3) We observed that as the number of iterations increases from 200 to 600, the apply rate decreases. This phenomenon indicates that as the number of iterations increases, the model may be affected by some interference information when generating results, resulting in a decrease in the quality of the generated results. Therefore, when selecting the number of iterations, it is necessary to consider avoiding the influence of excessive interference information.
4.5. Evaluation on In-house Dataset
Language Resolved Fault_Location
Java 14.7% 41.2%
TypeScript 18.8% 28.1%
JavaScript 17.2% 31.3%
Average 16.9% 33.5%
Table 7. Results on Multilingual In-house Dataset.
Tasks Automated Human-in-the-Loop
Resolved 16.7% 43.3%
Fault_Loc 40.0% 66.7%
Table 8. Results of Human-in-the-Loop Intervention on Alibaba Cloud In-house Dataset Subset.
To assess the effectiveness of LingmaAgent in real-world industrial settings, we conducted a comprehensive evaluation using an in-house dataset meticulously curated from Alibaba Cloud’s diverse development scenarios. This dataset was designed to test the performance of LingmaAgent on multi-language repositories, focusing on Java, JavaScript, and TypeScript - three of the most prevalent languages in cloud-based and web application contexts. The dataset encompasses a wide range of projects including e-commerce platforms, cloud infrastructure services, and data analytics tools, representing the complexity and diversity of industrial-scale software projects. It comprises 10 Java repositories (averaging 1538 files and 3.4 issues each), 24 JavaScript repositories (averaging 503 files and 4 issues each), and 16 TypeScript repositories (averaging 793 files and 4 issues each), for a total of 194 issues.
We deployed LingmaAgent with the GPT-4 for this evaluation. We used the same metrics to evaluate LingmaAgent’s performance. The experiment was conducted in two phases:
  • Fully Automated Resolution. LingmaAgent attempted to resolve issues without any human intervention. For this phase, we conducted a full evaluation on all 194 issues in the dataset.
  • Human-in-the-Loop Intervention. For issues not resolved in the first phase, we implemented a human-in-the-loop approach. Development engineers intervened in the product interaction pipeline, manually adjusting LingmaAgent’s generation plans and search_api call, and refining potential fault localizations (interventions must less than 5 times). The final patches were then generated using the model. For this phase, we randomly selected 30 issues for manual evaluation.
The results of our evaluation are presented in Table 7 and 8. These findings offer valuable insights into LingmaAgent’s performance in industrial-scale software engineering tasks. In the Fully Automated Resolution phase, LingmaAgent successfully resolved 16.9% of the issues across the multilingual dataset. This result shows that the system is capable of autonomously handling some real software engineering tasks without any human intervention, but there is still much room for improvement. For the Human-in-the-Loop Intervention phase, we randomly selected a subset of 30 issues. In this subset, LingmaAgent’s automated performance was consistent with the full dataset, resolving 16.7% of issues independently. However, with human intervention, the resolution rate dramatically increased to 43.3%. This substantial improvement of 26.7 percentage points highlights the synergistic potential of human-AI collaboration in tackling complex software engineering problems. This indicates that the system effectively augments human problem-solving skills rather than replacing them. Overall, these results demonstrate LingmaAgent’s potential as a powerful tool in industrial software engineering contexts, particularly when integrated into a collaborative workflow with human developers.
@@ 5. Limitation
Resource Overhead. Although LingmaAgent aims to guide LLMs to fully understand the whole software repository to effectively solve the challenges in ASE, the MCTS process does require a certain amount of resource consumption. Specifically, we set the maximum number of iterations to 600 and the maximum search time to 300 seconds to ensure that the model can fully explore the search space and accurately evaluate the rewards of different paths. However, such settings are controllable and adjustable to adapt to different application scenarios and resource constraints. Through reasonable parameter adjustment, the best balance between resource consumption and result accuracy can be found. In addition, as shown in Table 6, only 50 iterations can also achieve results that are superior to other agents. At the same time, in Table 2, we ran LingmaAgent on different base models. We found that with the introduction of the next generation of models, the improvement of model capability and the cost will also bring about the improvement of LingmaAgent effect and the reduction of cost, which further demonstrates the possibility of application with LingmaAgent. Further research may discover more efficient strategies to reduce resource requirements while maintaining or improving agents performance.
Evaluation of LingmaAgent. While our evaluation of LingmaAgent demonstrates promising results, several limitations in our current assessment approach warrant consideration. Our primary evaluation relies on the SWE-bench Lite dataset, and although we conducted additional tests using an in-house dataset from Alibaba Cloud, the scope of our human-in-the-loop evaluation was limited due to the substantial human resources required for comprehensive interaction and assessment. This constraint potentially limits the generalizability of our findings to a broader range of real-world scenarios. Additionally, there is a possibility that some of the models we used, may have been exposed to parts of the repositories in our test set during their training. While this potential data leakage is a concern, we believe that the relative performance improvements demonstrated by LingmaAgent still provide valuable insights into its effectiveness. Nevertheless, this limitation highlights the need for more controlled evaluation environments in future studies. To address these limitations and enhance the robustness of future evaluations, we propose several directions for future work: creating standardized protocols for human-in-the-loop evaluations, and developing a dynamic, continuously updated version of SWE-bench that evolves with the software engineering field. By addressing these limitations and expanding our evaluation methodologies, we aim to provide more robust and generalizable assessments of AI-assisted software engineering tools like LingmaAgent in the future.
@@ 6. Conclusion
This paper emphasizes the importance of understanding entire software repositories for achieving Automated Software Engineering. We introduce LingmaAgent, a novel LLM-based agent method that comprehensively analyzes repositories through knowledge graph construction, MCTS-enhanced exploration, and global experience-based planning. This approach enables agents to solve real-world GitHub issues effectively. Extensive experiments demonstrate LingmaAgent’s superior performance over existing systems on the SWE-bench Lite benchmark. Ablation studies highlight the significance of global repository experiences and the potential of integrating runtime feedback. We also validate LingmaAgent’s effectiveness in real-world industrial settings using an Alibaba Cloud dataset, showcasing its capabilities in both automated and human-in-the-loop scenarios.
Future work will focus on developing a dynamic, evolving version of SWE-bench, and optimizing resource efficiency while maintaining or improving agent performance. These efforts aim to advance the field of AI-assisted software engineering and provide more robust solutions for complex software development tasks.